a rectangle has a perimeter of 78 meters. the lenth is 11 meters longer than triple the with. find the dimensions of the rectangle.
Answer by Paul(988)

Width = x
Length = 3x+11
Formula 2(l+w)=P
2(x+3x+11)=78
8x+22=78
8x=56
x=7
3(7)+11=32
Hence, the length is 32m and the width is 7m.
Paul.